OpenStudy (anonymous):

Need help with a trig problem, will FAN and MEDAL. Problem will be in comments

OpenStudy (anonymous):

\[\sin \theta=\frac{ -4 }{ 5 } \ , \frac{ -\pi }{ 2 }<\theta<0\] \[\cos \beta=\frac{ -5 }{ 13 } , \frac{ \pi }{ 2 }<\beta<\pi\] What is cos theta/2? I tried it and I need to check my answer

OpenStudy (anonymous):

\[\huge \sqrt{\frac{1+\frac{3}{5}}{2}}\] i thinkor maybe it is negative

OpenStudy (anonymous):

\(\theta\) is in quadrant 4, making \(\cos(\theta)\) positive \[\sin(\theta)=-\frac{4}{5}\implies \cos(\theta)=\frac{3}{5}\]

OpenStudy (anonymous):

plug in the the half angle formula and you get what i wrote above, except it is \(\pm\) and we have to figure out whether it is positive or negative

OpenStudy (anonymous):

i guess positive, because if \[-\frac{\pi}{2}<\theta<0\] then also \[-\frac{\pi}{2}<\frac{\theta}{2}<0\]

OpenStudy (anonymous):

also you have some arithmetic to do, i will leave that up to you

OpenStudy (anonymous):

In the end I came up with \[2\sqrt{5} / 5\]